Fermi National Laboratory


Accelerator Update for December 16, 2005 - December 19, 2005

Friday December 16

The day shift began with the Recycler Ring (RR) stash at 186.6 E10, the Antiproton Source (Pbar) stack at 74.4 E10, and with MiniBooNE, and Meson MCenter taking beam.

Shot setup begins at 11:30 AM.

Operations established store 4553 at 2:55 PM with an initial luminosity of 98.3 E30.

Main Injector (MI) was turned off at 2:10 PM to look for NuMI LCW leak.

At 3:06 PM, Linac Klystron RF Station #6 (KRF6) trips off and will not reset. All Proton Source beam is off, experts are investigating.

NuMI LCW leak is found and MI is turning back on at 5:07 PM.

Saturday December 17

The midnight shift began with Operations monitoring store 4553, the Pbar stack at 133.1 E10, and with KRF6 repairs still on going.

KRF6 was repaired (the transformer replaced) at 3:00 AM.

Operations put the Recycler into shot setup at 3:19 AM to stash antiprotons.

Pbar returns to stacking at 4:47 AM, beam to MIPP.

The Recycler Ring is experiencing cooling problems, experts are investigating.

Store 4553 was terminated at 4:02 PM.

Shot setup begins at 4:38 PM.

Operations established store 4554 at 7:08 PM with an initial luminosity of 87.01 E30.

At 7:19 PM the MI/RR are being prepared for access to investigate the Recycler Ring cooling problems.

Recycler cooling work was completed by 9:05 PM, the MI and Recycler are being turned back on.

Beam to MiniBoone at 9:37 PM. Stacking and beam to MIPP is being held off due to a problem with a power supply (I:HVF12).

Sunday December 18

The midnight shift began with Operations monitoring store 4554, the Pbar stack at 26.7 E10, MiniBoone taking beam, and experts working on I:HVF12 .

Repairs to I:HVF12 were completed at 2:09 AM, Pbar returned to stacking, beam also resumed to MIPP.

Operations put the Recycler into shot setup at 10:05 AM to stash antiprotons.

Pbar returns to stacking at 11:03 AM, beam returns to MIPP.

Lost the store at 2:02 PM due to a Tevatron quench at F4, the quench was caused by a separator spark at D1. Recovery is underway.

Shot setup begins at 4:49 PM.

Operations established store 4555 at 7:31 PM with an initial luminosity of 93.06 E30.

Monday December 19

The midnight shift began with Operations monitoring store 4555, the Recycler Ring (RR) still experiencing cooling problems, Pbar stacking, and with MiniBooNE, and MIPP taking beam.

The Plan for Monday

Keep the store, access the Main Injector to fix Recycler cooling problems, run beam to MiniBoone, stack, stash and MCenter. Bring NuMI back on line this afternoon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/19/2005   email Fermilab